Question 50 of 75 Tyler purchased a rental house in May 2011 for $135,000, of which $35,000 was attributed to the land. Tyler sold the renta house in August 2017. How much will he deduct for depreciation in 2017? 1O $1.364 $2,273 $2,425 $3,636 Mark for follow up « Back Save / Return Later Summary Next »
Answer #1
Q 50)
Answer: $2425
=> As per MACRS depreciation system, the rental property
the recovery period is 27.5 years, so
In the problem, The building value is $135000 - $35000 = $100000
The depreciation recovery amount per year = $100000/27.5 years = $3636.36
The depreciation for the year 2017 is for 8 months.
The depreciation for the year 2017= 3636.36*8/12 = $2425